### v1

@@ 3.4 Research Plant with MPC
In this case study, we identify a hybrid model for a research plant and embed it as a controller model in a MPC framework. The flowsheet of the research plant is shown in Fig. 8. The plant comprises a large water reservoir and three water buffer tanks. A pump circulates the water in the pipeline, equipped with a pressure controller (PIC 1). The three buffer tanks are hydraulically connected, allowing water to flow between them and the reservoir. Tanks 1 and 3 are connected to the circulation pipeline, and their feed flow rates are controlled by flowrate controllers (FIC 1 and FIC 2). By setting flowrate setpoints for FIC 1 and FIC 2, Tanks 1 and 3 can be filled with water. Each buffer tank is equipped with level indicators (LI 1, LI 2, and LI 3), and the reservoir holdup is measured by LI 4. The flow rates between the buffer tanks and from Tank 3 to the reservoir are not directly measurable.
@@
with the holdups $h_{1}$, $h_{2}$, $h_{3}$ of Tanks 1, 2, 3, respectively, $h_{\mathrm{res}}$ the holdup of the water reservoir, the input flowrates $F^{\mathrm{in}}1$ and $F^{\mathrm{in}}2$ in Tank 1 and 2, respectively, and the streams $F{i,j}$ from Tank $i$ to $j$, and $F\mathrm{3,r}$ from Tank 3 to the reservoir. We consider the streams $F_{i,j}$ and $F_{\mathrm{3,r}}$ unknown, i.e., we model these variables with data-driven model parts using the incremental identification approach. We identify a hybrid model based on the following formulation:
@@
We use the identified hybrid model as a controller model in MPC to control the research plant. Fig. 10 shows the closed-loop results, demonstrating that the MPC with the embedded hybrid model successfully tracks the desired setpoints. Note that stationary offsets can be addressed by offset-free NMPC [32, 33], which is outside the scope of this work.
4 Conclusions
@@
The above approach offers several significant advantages over traditional simultaneous identification methods. The separation of mechanistic and data-driven components allows independent evaluation and optimization of each model part, significantly simplifying the identification process and enabling more efficient model development. It also allows for early detection and isolation of modeling issues, whether they originate from mechanistic or data-driven components. By avoiding the embedding of machine learning models within complex dynamic optimizations, our approach enables the use of standard machine learning techniques and libraries. Furthermore, the methodology’s regularization approach enables the identification of suitable dynamic hybrid models even with limited measurement data, making it particularly valuable for industrial applications.
@@
As already mentioned, the effectiveness of the proposed approach relies on a careful choice of the regularization parameters $\boldsymbol{W}_{k}^{R}$ (cf. Equation (3)). Future work could focus on the development of systematic approaches for selecting regularization parameters, something which could further enhance the method’s applicability in industrial settings.

### v2

@@ 3.4 Research Plant with MPC
In this case study, we identify a hybrid model for a research plant and integrate it within an MPC framework for process control. As shown in Figure 9, the plant comprises a large water reservoir and three water buffer tanks. A pump circulates the water in the pipeline, equipped with a pressure controller (PIC 1). The three buffer tanks are hydraulically connected, allowing water to flow between them and the reservoir. Tanks 1 and 3 are connected to the circulation pipeline, and their feed flow rates are controlled by flowrate controllers (FIC 1 and FIC 2). By setting flowrate setpoints for FIC 1 and FIC 2, Tanks 1 and 3 can be filled with water. Each buffer tank is equipped with level indicators (LI 1, LI 2, and LI 3), and the reservoir holdup is measured by LI 4. The flow rates between the buffer tanks and from Tank 3 to the reservoir are not directly measurable.
@@
with the holdups $h_{1}$, $h_{2}$, $h_{3}$ of Tanks 1, 2, 3, respectively, $h_{\mathrm{res}}$ the holdup of the water reservoir, the input flowrates $F^{\mathrm{in}}1$ and $F^{\mathrm{in}}2$ in Tank 1 and 2, respectively, and the streams Fi,jF{i,j} from Tank $i$ to $j$, and $F\mathrm{3,r}$ from Tank 3 to the reservoir. We consider the streams $F_{i,j}$ and $F_{\mathrm{3,r}}$ unknown, i.e., we model these variables with data-driven model parts using the incremental identification approach. We identify a hybrid model based on the following formulation:
@@
We use the identified hybrid model within the MPC to control the research plant. Fig. 11 shows the closed-loop results, demonstrating that the MPC with the embedded hybrid model successfully tracks the desired setpoints. Note that stationary offsets can be addressed by offset-free NMPC [34, 24], which is outside the scope of this work.
4 Concluding Remarks
@@
As described in the Introduction (Section 1) of this paper, the above approach offers several significant advantages over traditional simultaneous identification methods. It is worth noting that numerical comparisons of the two approaches have been reported in the literature [15] for particular classes of models. Our attempts to perform similar comparisons for the general dynamic hybrid systems considered in this paper were less successful, with the simultaneous approach facing significant numerical difficulties. There appeared to be two main causes for these failures:
  • Overparameterization: For problems with limited experimental data (cf. Case Study 2 in Section 3.3), the estimation problem is vastly overparameterized because of the large number of parameters introduced by even relatively small neural networks. The incremental approach overcomes this problem via regularization, but there is no easily implementable way of achieving an equivalent effect in the simultaneous approach.
  • Numerical convergence: as the neural network parameters have no physical meaning, the parameter estimation invariably starts from very poor initial guesses, which often results in physically unrealistic time profiles for the quantities $\boldsymbol{p}$. This problem is common to both the simultaneous and the incremental approaches. However, in the former, $\boldsymbol{p}$ is embedded within an (often complex) set of nonlinear differential and algebraic equations (DAEs); therefore, having a completely unphysical behavior for $\boldsymbol{p}$ during the early stages of the parameter estimation can create major numerical problems (e.g. with the stable integration of the system). In contrast, the incremental approach simply operates on a table of input/output data, and uses standard machine learning algorithms and tools which are well suited for dealing with bad initial parameter estimates.
@@
The effectiveness of the proposed approach relies on a careful choice of the regularization parameters $\boldsymbol{W}_{k}^{R}$ (cf. Equation (3)). As described in Section 3.2.6, in this paper we adopted an empirical, trial-and-error approach for this task. Future work could focus on the development of more systematic approaches, something which could further enhance the method’s applicability in industrial settings.
Finally, we note that our paper assumes that the structure of the hybrid model to be identified is already chosen by the model’s developer. Consequently, the methodology presented is based on a formulation (Equation (1)) that is general enough to accommodate a wide range of such structures. One of the simplest such structures is obtained by the parameters $\boldsymbol{p}$ being added as residual terms on the right-hand sides of the differential equations (e.g. mass, energy or momentum balances). This approach can work well for models involving a small number of states, such as those considered in our first and third examples (Sections 3.2 and 3.4, respectively). However, many models of practical interest to process engineering applications involve hundreds or thousands of differential states. In such cases, introducing an equal number of residual terms and attempting to fit a neural network which expresses them as functions of the other model variables could be problematic, particularly from the point of view of identifiability given the often limited availability of experimental data. Instead, the model’s developer may wish to introduce a much smaller number of parameters to describe specific physical phenomena that are not adequately described by mechanistic equations or correlations. Obviously, a degree of physical insight is required to be able to correctly identify these gaps in physical understanding. The development of techniques for identifying such gaps automatically would indeed be a worthwhile topic for further research. Systematic approaches for model structure determination for tailored system-theoretical properties, e.g., based on semi-infinite programming as in [24], may be relevant in this context.
Acknowledgments
The authors acknowledge the use of Microsoft Copilot  [^9] Microsoft Copilot, https://learn.microsoft.com/en-us/copilot/, accessed: 10.08.2025, an AI-powered writing assistant, in the drafting and refinement of this manuscript.